Shahryari

1 ℤ_{2}-graded symmetry classes of tensors

By M. Shahryari.

In this paper, we define a natural ℤ_{2}-gradation on the symmetry class of tensors V_{χ}(G). We give the dimensions of even and odd parts of this gradation. Also we prove that the even part ( the odd part) of this gradation is zero, if and only if the whole symmetry class is zero.